Mission Operations Days: 2004/101 to 2004/107
April 10, 2004 00:00:00 GMT to April 16, 2004 00:00:00 GMT
Terra Spacecraft and MODIS Instrument Status:
|AM-1 "Terra" is in Normal Mode|
PFM MODIS is in A-Side Nominal Science Mode
(Cross Strapped with FR B)
|Blackbody||A On (290K); B Off||Nominal Science Mode|
|Calibration Electronics||A On; B Off||Nominal Science Mode|
|Control Processor||A On; B Off||Nominal Science Mode|
|Nadir Door||Unlatched, Open||Nominal Science Mode|
|Space View Door||Unlatched, Open||Nominal Science Mode|
|Solar Diffuser Door||Unlatched, Screened||Permanent Configuration|
|FDDI Formatter||A On; B Off||Nominal Science Mode|
|FDDI Port||A On; B Off||Nominal Science Mode|
|FIFO Memory||#1 & #2 On; #3 & #4 Off||Nominal Science Mode|
|Format Processor||A Off; B On||Nominal Science Mode|
|Power Supplies||#1 On, #2 Off||Nominal Science Mode|
|PV VIS FPA||A On; B Off||Nominal Science Mode|
|PV NIR FPA||A On; B Off||Nominal Science Mode|
|PV SMIR FPA||A On; B Off||Nominal Science Mode|
|PV LWIR FPA||A On; B Off||Nominal Science Mode|
|PC LWIR FPA||A On; B Off||Nominal Science Mode|
|RC Outgas Heaters||All 3 (CS, IS, OS) Off||Nominal Science Mode|
|RC LWIR FPA Heater||On (83K)||Nominal Science Mode|
|RC SMIR FPA Heater||Off||Nominal Science Mode|
|Scan Assembly||A On; B Off||Nominal Science Mode|
|SDSM||Off||Nominal Science Mode|
|SRCA||Off||Nominal Science Mode|
|PS1 Survival Heater||Enabled||Nominal Science Mode|
|PS2 Survival Heater||Enabled||Nominal Science Mode|
|Timing Generator||A On; B Off||Nominal Science Mode|
|Flight Software||Rev BD + 11 sets of patches||Nominal Science Mode|
|Inhibit Ids Set||none||Nominal Science Mode|
|TMON 66||Disabled||Nominal Science Mode|
|TMON 67||Disabled||Nominal Science Mode|
|TMON 68||Enabled||Nominal Science Mode|
|TMON 69||Disabled||Nominal Science Mode|
This Week’s Completed MODIS Activities:
Saturday, April 10, 2004
Sunday, April 11, 2004
Monday, April 12, 2004
Tuesday, April 13, 2004
2004/104 20:29:00-20:37:48 ATC MOD_OA16_SDSM_SCRN_NO_DOOR_A.4 OA-16: SD/SDSM screened side A, no door movement.
2004/104 21:10:00 ATC MOD_SRCA_NOMINAL_SETUP.4 Setup of nominal lamp configuration for SRCA calibration
2004/104 21:18:01-21:53:11 ATC MOD_OA19_SRCA_FULL_RAD_A.6 OA-19: SRCA Full Radiometric Calibration.
Data collection from 21:33:55-21:53:10.
Wednesday, April 14, 2004
Thursday, April 15, 2004
Friday, April 16, 2004
2004/107 22:00:00 ATC MOD_SRCA_NOMINAL_SETUP Nominal setup activity for SRCA activities
2004/107 22:04:26-22:39:33 ATC MOD_SPECIAL_10W_10WND_1W SRCA radiometric mode calibration with 10W, 10W + ND filter, and 1W lamp levels
Data collection from 22:05:30-22:38:30
This Week’s Scheduled MODIS Activities Not Completed:
Upcoming MODIS Events:
Calibrations for April will include weekly SDSM, monthly SRCA 10W Radiometric, bimonthly PV Ecal, and quarterly SRCA Full Spectral and Black Body calibrations.
Saturday, April 17, 2004
2004/108 21:00:00 ATC MOD_SRCA_BACKUP_SETUP Setup activity for SRCA activities using backup lamp #4 for 10W lamp level and 1W lamp #2 for the 1W lamp level.
2004/108 21:08:44-21:43:51 ATC MOD_SPECIAL_10W_10WND_1W SRCA radiometric mode calibration with 10W, 10W + ND filter, and 1W lamp levels
Data collection from 21:09:48-21:42:48
Tuesday, April 20, 1994
2004/111 20:34:35-20:43:23 ATC MOD_OA16_SDSM_SCRN_NO_DOOR_A.4 OA16: SD/SDSM screened side A, no door movement.
2004/111 22:22:29-22:25:00 ATC MOD_OA27_PV_ECAL.3
Wednesday, April 21, 2004
2004/112 00:30:00 ATC MOD_SRCA_NOMINAL_SETUP.4 Sets up nominal A Side SRCA lamp configuration
2004/112 00:38:37-01:30:18 ATC MOD_OA22_SRCA_FULL_SPECT_1A.6 OA-22: SRCA Full Spectral Calibration Part 1 - A Side
30W lamp data collection is from 00:54:31-01:30:17 (in S/C night)
2004/112 02:34:22-03:17:12 ATC MOD_OA22_SRCA_FULL_SPECT_2A.5 OA-22: SRCA Full Spectral Calibration Part 2 - A Side
30W lamp data collection is from 02:34:22-03:08:52 (in S/C night)
2004/112 03:50:00 ATC MOD_SRCA_NOMINAL_SETUP.4 Sets up nominal A Side SRCA lamp configuration
2004/112 03:56:24-04:47:42 ATC MOD_OA22_SRCA_FULL_SPECT_3A.6 OA-22: SRCA Full Spectral Calibration Part 3 - A Side
10W lamp data collection is from 04:12:51-04:47:24 (in S/C night)
2004/112 05:47:24-06:32:41 ATC MOD_OA22_SRCA_FULL_SPECT_4A.4 OA-22: SRCA Full Spectral Calibration Part 4 - A Side
10W lamp data collection is from 05:47:24-06:25:24 (in S/C night)
Friday, April 23, 2004
2004/114 02:00:00 ATC MOD_OA26_BB_270K_A.1 Set blackbody temperature to 270K on the A Side
2004/114 20:00:00 ATC MOD_OA26_BB_280K_A.1 Set blackbody temperature to 280K on the A Side
There will be no Lunar Calibration Rolls in April, May, or June for the Terra MODIS. The necessary roll angles are -26.3051, -27.2921, and -22.1398 degrees, respectively; all of which are beyond the -20.0 degree limit.
Terra performed a Drag Make Up (DMU) Maneuver on Tuesday, April 6, 2004 (DOY 2004/097). Time of the maneuver was 18:30:34 GMT and lasted for 20 seconds. Loss of, and regaining of, high fidelity pointing was 17:00:34 GMT and 21:00:34 GMT, respectively.
MODIS NAD (Nadir Aperture Door) Anomaly:
Telemetry still indicates that the “NAD Open” (MOD_CR_DR_NAD_OPEN) trigger is set to the “NOT_OPEN” position; however, science data has shown that the NAD is open. The NAD OPEN trigger may yet “pop” to OPEN due to thermal cycling on the trigger, as was seen on 2002/079.
General Instrument Comments:
MODIS is in Science Mode, using all A-side components, but cross-strapped to Formatter B, with the blackbody maintained at 290K and the Solar Diffuser Door at the Screened position.
MODIS Telemetry Trends:
Nominal telemetry trending is ongoing.
Non-MODIS Significant Events:
Terra Solid State Recorder Anomaly
The SSR Anomaly is still under investigation. MODIS currently is operating with 34 Supersets allotted for its buffer. There are plans in the works for possibly power cycling the SSR to get back PWA’s that are in the Bad List, but no date has yet been set. There are no current plans for a change in science data capture rates. For details, see previous reports.
The EMOS mission status webpage (which includes instrument data loss records) can be found at the following URL. Note that a password is now required to access the EOS mission links web pages. There are directions at the URL below for how to obtain a username/password.
Limited Life Item Status:
SRCA 10W Lamp #1: 227.5 of 500 hours
SRCA 10W Lamp #2: 166.2 of 500 hours
SRCA 10W Lamp #3: 175.4 of 500 hours
SRCA 10W Lamp #4: 62.2 of 500 hours
SRCA 1W Lamp #1: 565.9 of 4000 hours
SRCA 1W Lamp #2: 276.7 of 4000 hours
Solar Diffuser Door: 2146 of 3022 Movements
Nadir Aperture Door: 538 of 1316 Movements
Space View Door: 441 of 1316 Movements